OpenStudy (kenialove):

The relation that is represented by the equation y = 45 - 3x has the domain { 5,7,9,11 } . Which of the following sets is the range of the relation ?

OpenStudy (saurabh135):

{ 30, 24, 18, 12}

OpenStudy (kenialove):

How is that the answer thou ?

OpenStudy (zzr0ck3r):

put in 5, get out 30 put in 7, get out 24 put in 9, get out...

OpenStudy (kenialove):

I still don't understand .. Wat do u mean put in and get out ?

Vocaloid (vocaloid):

the domain {5,7,9,11} tells you what x-values to use and the equation y = 45-3x tells you what y values to get when we "put in" x = 5, we get "y=30" as our answer, so 30 becomes part of the range keep going for all the x-values

OpenStudy (anonymous):



OpenStudy (anonymous):

they put these silly pictures of "function machines" in modern textbooks, thought i would try my hand at one

OpenStudy (kenialove):

But how does 5 give u 30 and 7 give u 24 and 9 give u 18 and 11 give u 12 ?

Vocaloid (vocaloid):

let's "put in" x = 5 x = 5 y = 45 - 3x = ?

OpenStudy (kenialove):

Oohhhhh I get it thanks so much .. so it's like this 45 - 3x = y 45 - 3(5) = y 45 - 15 = 30